OpenStudy (anonymous):

Part 1: [4 points] Find the measures of central tendency (mean, median, mode and range) for the following data set: 14, 27, 24, 29, 28, 25, 14, 26 Part 2: [3 points] Tell which of these measures would best represent the whole data set and why.

OpenStudy (anonymous):

1. get them in numerical order first

OpenStudy (anonymous):

14,14,24,25,26, 27,28,29

OpenStudy (anonymous):

Mean= average. add together all the values, and divide by the number of values. 14+14+24...=187

OpenStudy (anonymous):

since there are 8 numbers, divide 187 by 8 23.375 mean=23.375

OpenStudy (anonymous):

ur left with 25,26 in the middle. find the average of those two. 25+26/2=25.5 median=25.5

OpenStudy (anonymous):

mode=most. which number occurs the most times in this set? answer=14 mode=14

OpenStudy (anonymous):

Range= biggest number in ur set minus the smallest number 29-14 15 Range=15

OpenStudy (anonymous):

2. mean=23.375 median=25.5 mode=14 range=15

OpenStudy (anonymous):

@dan815 need help last part

OpenStudy (anonymous):

@Hannnnah almost done

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

adios! aie aie aie!

OpenStudy (dan815):

hmmm what would best represent the whole data set id have to say Range, or Mean, but probably range since it will tell you what range of data we are dealing with, mean is not bad to go with too

OpenStudy (anonymous):

oh ok thx

OpenStudy (dan815):

The range should say 14-29